# Instructions for *ACL Proceedings

### v1

@@ Abstract
LLM-as-a-Judge refers to the automatic modeling of preferences for responses generated by Large Language Models (LLMs), which is of significant importance for both LLM evaluation and reward modeling. Although generative LLMs have made substantial progress in various tasks, their performance as LLM-Judge still falls short of expectations. In this work, we propose Think-J, which improves generative LLM-as-a-Judge by learning how to think. We first utilized a small amount of curated data to develop the model with initial judgment thinking capabilities. Subsequently, we optimize the judgment thinking traces based on reinforcement learning (RL). We propose two methods for judgment thinking optimization, based on offline and online RL, respectively. The offline method requires training a critic model to construct positive and negative examples for learning. The online method defines rule-based reward as feedback for optimization. Experimental results showed that our approach can significantly enhance the evaluation capability of generative LLM-Judge, surpassing both generative and classifier-based LLM-Judge without requiring extra human annotations [^1] Codes and data are openly available at https://github.com/HuihuiChyan/Think-J.
Think-J: Learning to Think for Generative LLM-as-a-Judge
[…]
footnotetext: * The first three authors contributed equally.
@@ 1 Introduction
As the capabilities of generative LLMs continue to advance, accurately evaluating the response quality has emerged as a crucial challenge. This is not only vital for more efficient model development and comparison but also essential in the context of Reinforcement Learning from Human Feedback (RLHF), which relies on precise preference modeling as guidance Wang et al. (2024a). However, traditional evaluation methods for generative models, such as BLEU Papineni et al. (2002), are based on predefined reference answers, which are often unavailable in open-ended scenarios.
@@
Some studies have proposed LLM-as-a-Judge Zheng et al. (2023), which leverages the generative capabilities of LLMs for evaluating response quality. These work either directly leverage proprietary LLMs or fine-tune a smaller judge based on preference data Gu et al. (2024). However, the accuracy of their judgments remains unsatisfactory as revealed by recent benchmarks Lambert et al. (2024). Other research has suggested fine-tuning a classifier based on preference data Liu et al. (2024a). While this method can achieve higher judgment accuracy, it lacks interpretability due to its scalar output, and the performance is highly dependent on the data quality Wang et al. (2024b).
Inspired by recent reasoning models such as o1 Jaech et al. (2024) and Deepseek-R1 Guo et al. (2025), in this work, we propose Thinking-enhanced Generative Judge (Think-J). Think-J aims to train a better generative judge by optimizing the model’s judgment thinking capabilities. Specifically, our method consists of two steps:
@@
2) Judgment Thinking Optimization. Due to the lack of high-quality critique annotation in preference datasets, we opt to optimize judgment thinking ability based on reinforcement learning (RL). Specifically, we adopted two methods: a) Critic-guide Offline Learning, leveraging an additional critic model to generate corresponding thinking traces based on provided judgment results, thus constructing positive and negative examples for offline RL. b) Rule-based Online Learning, defining rule-based rewards based on the correctness of judgment results, thus optimizing the thinking trace by online RL Shao et al. (2024).
@@
    We propose to stimulate the judgment thinking ability of generative models with a small amount of carefully curated data.
@@
    We propose to optimize the judgment thinking ability of generative models with the correctness of judgment results as feedback.
@@
    Our proposed Think-J significantly outperforms previous generative and classifier-based LLM-judges for preference modeling.
2 Related Work
After the emergence of LLMs, numerous efforts have been made to design a more effective method for LLM evaluation Chang et al. (2023). One of the most scalable and effective method is LLM-as-a-Judge Li et al. (2023b); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT4 Achiam et al. (2023), to evaluate the LLM’s response. For example, AlpacaEval Li et al. (2023b) used the win rate compared with baseline response determined by GPT-4 as the evaluation result. MT-Bench Zheng et al. (2023) automatically scored the model’s answers using GPT-4 as the results. The GPT-4-based evaluator is proven to presents comparable or even better consistency and accuracy compared with human.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, follow-up works suggest fine-tuning language models locally for evaluations, including JudgeLM Zhu et al. (2023b), Auto-J Li et al. (2023a), Prometheus Kim et al. (2023), Prometheus-2 Zhu et al. (2023a), OffsetBias Park et al. (2024), etc. These work typically construct preference data with judgment annotations and then finetune open-sourced LLMs to generate the judgment. Despite these fine-tuned judge models all achieve comparable accuracy with proprietary models, the evaluation is mostly conducted on the in-domain testsets, and these works are verified with a low scalability on more general benchmarks Huang et al. (2024).
Another group of work fine-tunes a classifier on preference data based on the Bradley-Terry model, which is more commonly used on reward modeling. This approach is simple yet effective, as demonstrated on RewardBench Lambert et al. (2024) where most top-performing models are trained in a classification style Liu et al. (2024a). However, this method does not fully leverage the generative capabilities of LLMs, and is unable to provide rationales for its judgments, which is crucial for scalable evaluation. While recent work has begun to leverage the generative abilities of LLMs to combine critiques for scalar reward prediction Ke et al. (2024); Ye et al. (2024), these critiques are often distilled from stronger proprietary models and hardly influence the final prediction Liu et al. (2025b). Effectively integrating the generative abilities of LLMs into evaluation remains an open challenge Chen et al. (2025a); Whitehouse et al. (2025); Chen et al. (2025b); Wang et al. (2025a).
3 Approach
@@ 3.1 Judgment Thinking Initialization
Recent studies have shown that LLMs inherently possess long chain-of-thought (CoT) reasoning capabilities, which can be activated with a small amount of data Muennighoff et al. (2025); Ye et al. (2025). In this work, we also curate high-quality preference data, LIMJ707, to initialize the thinking capability of the judge model. LIMJ707 are selected based on three principles:
  • Accuracy: The judgment annotation should be correct. We leverage the high-quality preference data Skywork-Preference-v0.2222huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.2, which has been carefully validated to ensure accuracy.
@@
  • Diversity: The instruction should encompass various types to enhance judgment thinking capabilities in different aspects. We represent the instructions with an embedding model and then merge samples that are too close to each other [^3] For more details please refer to Appendix B.1..
The data statistics during constructing LIMJ707 are shown in Table 1. Based on LIMJ707, we construct judgment thinking trace by Deepseek-R1. After that, the annotated samples are used to initialize the model with judgment thinking capability by Supervised Fine-tuning Ouyang et al. (2022).
@@
Despite its superior reasoning capability, the thinking trace generated by R1 are excessively long and could possibly cause difficulty for further optimization and increase computational overhead. Specifically, for our case, we hypothesize that general LLM judgment does not necessarily require a detailed, step-by-step reasoning process. On the contrary, identifying one key factor is sufficient to make correct judgment in most cases.
Therefore, we introduce Trace Clipping to enhance the effectiveness of thinking initialization [^4] An example is presented in Appendix B.1.. Specifically, we observed that the CoTs generated by R1 are composed of two parts: a lengthy reasoning process, and an explanation which summarizes the reasoning process. Therefore, we remove the first part and use the second part as the output trace. This not only facilitates subsequent optimization but also enhances the judgment readability.
@@ 3.2 Judgment Thinking Optimization
To further enhance the alignment between the judge and human preference, the initialized thinking trace should be further optimized on preference data. However, preference data typically only includes binary labels without thinking trace annotations. Therefore, we propose judgment thinking optimization based on reinforcement learning (RL). Specifically, we propose two optimization approaches based on offline and online RL methods, respectively, as detailed in the following sections.
3.2.1 Critic-guided Offline Learning
Offline RL methods represented by Direct Preference Optimization (DPO) Rafailov et al. (2023) has been widely applied to LLM pipelines due to their efficiency and simplicity Grattafiori et al. (2024). In this work, we also aim to optimize the thinking ability based on offline learning.
Due to the lack of golden thinking annotation in preference dataset, we propose to train an additional critic model [^5] Notice the critic model here differs from the critic model in traditional RL which is used for advantage estimation., to help to construct training samples. Both the critic and the judge models are trained on the same data (i.e., LIMJ707), with the following distinctions:
@@ 3.2.1 Critic-guided Offline Learning
With the help of critic model, samples with thinking annotation are constructed based on the correctness of judgment result. Therefore, the judge model will be enhanced to generate more accurate thinking, leading to more accurate judgment.
3.2.2 Rule-based Online Learning
The recent success of R1-style methods have demonstrated the effectiveness of online RL using discrete, rule-based rewards Shao et al. (2024). In this work, we also apply online rule-based RL approach to optimize the judgment thinking capability. More specifically, we mainly utilize the GRPO algorithm, whose optimization objective is:
@@ 3.2.2 Rule-based Online Learning
$\>\>\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(y|x)}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)}A_{i},\right.\right.$
@@
$r_{\text{accuracy}}$ =$=\begin{cases}1,&\text{if }\text{judgement}=\text{label}\\ 0,&\text{if }\text{judgement}\neq\text{label}\end{cases}$ (3)
$r_{\text{format}}$ =$=\begin{cases}0,&\text{if }\text{format is right}\\ -0.5,&\text{if }\text{format is wrong}\end{cases}$ (4)
$r_{\text{strength}}$ $=||\text{s}_{\text{pred}}-\text{s}_{\text{golden}}||,s\in\{1,2,3\}$ (5)
@@
Notice we incorporate a reward for assessing preference strength, which is defined as the degree to which the judge favors one response over another [^7] For example, a strength of 1 means the chosen response is only slightly better than the rejected, while a strength of 3 means the chosen is much better than the rejected.. The final judgment is also adjusted as:
@@
Preference strength helps to perceive the relative quality of response pairs, without uniformly providing the same reward for different pairs despite the quality gap. We employ a comparatively simple scale of reward scores of reward, as during actual training, we observed that the model tends to manipulate scores towards extreme values  [^8] For more details please refer to Appendix A.1..
While annotating an absolute score for a given response is challenging, annotating preference strength is relatively easy, making strength annotation more readily accessible Wang et al. (2024d, 2025b). On the other hand, if absolute score annotation is absent in the preference data, we can firstly fine-tune a BT-classifier based on the data, then leverage the classifier to assess the relative strength between chosen and rejected responses, as explained in Wang et al. (2024b).
Model Sample Num RewardBench RMBench Think-J
@@
Claude-3-5-Sonnet-20240620999www.anthropic.com/news/claude-3-5-sonnet 96.4 74.0 81.6 84.7 84.2 68.9 70.7
@@
GPT-4o-2024-08-06101010platform.openai.com/docs/models/gpt-4o 96.1 76.1 88.1 86.6 86.7 68.8 69.8
Gemini-1.5-Pro-0514111111deepmind.google/technologies/gemini/pro 92.3 80.6 87.9 92.0 88.2 74.4 68.1
JudgeLM-33B Zhu et al. (2023b) 100K 90.1 51.0 85.7 39.7 66.6 49.6 45.3
Prometheus-7b-v2.0 Zhu et al. (2023a) 40K 83.9 49.2 72.8 72.0 69.5 52.4 63.1
Prometheus-8x7b-v2.0 Zhu et al. (2023a) 40K 93.0 47.1 80.5 77.4 74.5 57.4 68.5
Llama-3-OffsetBias-8B Park et al. (2024) 276K 92.5 80.3 86.8 76.4 84.0 66.0 68.7
CompassJudger-32B Cao et al. (2024) 2041K 97.4 65.6 85.1 87.1 83.8 69.4 80.7
STE-Llama3.1-70B Wang et al. (2024c) 20K 96.9 85.1 89.6 88.4 90.0 65.3 72.0
SynRM-Command-R-35B Ye et al. (2024) 5K 97.5 76.8 88.5 86.3 87.3
CLoud-Llama3-70B Ke et al. (2024) 350K 98.0 75.6 87.6 89.0 87.6
Think-J-Qwen-2.5-32B 9.8K 96.7 83.2 90.1 92.0 90.5 79.8 75.8
Table 2: Experiment results of different LLM judges on RewardBench, RMBench and Auto-J.
Model Method HH-RLHF HelpSteer2-Preference
@@
Table 3: Experiment results of different LLM judge training methods on RewardBench.
@@ 4.1 Set-up
We mainly conducted experiments on two popular open-sourced models with their instruction version: Qwen-2.5 Qwen et al. (2025) and Llama-3 Grattafiori et al. (2024).
The optimization is conducted on two datasets: HelpSteer2-Preference121212huggingface.co/datasets/nvidia/HelpSteer2 and HH-RLHF131313huggingface.co/datasets/Anthropic/hh-rlhf. Notice while there are larger preference datasets available, we did not include them in our training, as our primary objective was to explore the most effective method for training LLM judges.
We compare Think-J with the following generative judgment approaches:
  • Direct Prompt Leverage the LLM to directly generate judgment.
  • SFT (w/o CoT) Train a generative model based on Supervised Fine-tuning to perform judgment without generating CoTs.
  • SFT (w/ CoT) Train a generative model on correct judgment thinking traces and results, which can also be deemed as Rejection Sampling.
We also compare Think-J with the following classifier-based approaches:
  • BT Classifier Feed the instruction and responses into the model and added a classification head, training it according to Bradley-Terry model.
  • CLoud Ke et al. (2024) First train the model to generate critiques, and then leverage the critiques as additional input to improve the BT Classifier.
  • SynRM Ye et al. (2024) Leverage critiques generated by a proprietary model as additional input to improve the BT Classifier. We use Deepseek-R1 to generate the critiques.
@@
We mainly perform evaluation on RewardBench Lambert et al. (2024), which is the most popular benchmark for evaluating preference modeling ability. We also evaluate our model on RMBench Liu et al. (2025a) and Auto-J-test Li et al. (2023a).
We report the best results achieved by either offline or online learning in the experiment results by default. We also mixed LIMJ707 into all training sets to ensure a fair comparison.

### v2

# Think-J: Learning to Think for Generative LLM-as-a-Judge

@@ Abstract
LLM-as-a-Judge refers to the automatic modeling of preferences for responses generated by Large Language Models (LLMs), which is of significant importance for both LLM evaluation and reward modeling. Although generative LLMs have made substantial progress in various tasks, their performance as LLM-Judge still falls short of expectations. In this work, we propose Think-J, which improves generative LLM-as-a-Judge by learning how to think. We first utilized a small amount of curated data to develop the model with initial judgment thinking capabilities. Subsequently, we optimize the judgment thinking traces based on reinforcement learning (RL). We propose two methods for judgment thinking optimization, based on offline and online RL, respectively. The offline method requires training a critic model to construct positive and negative examples for learning. The online method defines rule-based reward as feedback for optimization. Experimental results showed that our approach can significantly enhance the evaluation capability of generative LLM-Judge, surpassing both generative and classifier-based LLM-Judge without requiring extra human annotations.
footnotetext: * Equal contribution.footnotetext: $\dagger$ Corresponding Author.
Codehttps://github.com/huihuichyan/think-j
@@ 1 Introduction
As the capabilities of generative LLMs continue to advance, accurately evaluating the response quality has emerged as a crucial challenge(Huang et al. 2025). This is not only vital for more efficient model development and comparison but also essential in the context of Reinforcement Learning from Human Feedback (RLHF), which relies on precise preference modeling as guidance (Wang et al. 2024a; He et al. 2025; Liu et al. 2025c). However, traditional evaluation methods for generative models, such as BLEU (Papineni et al. 2002), are based on predefined reference answers, which are often unavailable in open-ended scenarios.
@@
Some studies have proposed LLM-as-a-Judge (Zheng et al. 2023), which leverages the generative capabilities of LLMs for evaluating response quality. These work either directly leverage proprietary LLMs or fine-tune a smaller judge based on preference data (Gu et al. 2024). However, the accuracy of their judgments remains unsatisfactory as revealed by recent benchmarks (Lambert et al. 2024). Other research has suggested fine-tuning a classifier based on preference data (Liu et al. 2024a). While this method can achieve higher judgment accuracy, it lacks interpretability due to its scalar output, and the performance is highly dependent on the data quality (Wang et al. 2024b).
Inspired by recent reasoning models such as o1 (Jaech et al. 2024) and Deepseek-R1 (Guo et al. 2025), in this work, we propose Thinking-enhanced Generative Judge (Think-J), as shown in Figure 1. Think-J aims to train a better generative judge by optimizing the model’s judgment thinking capabilities. Specifically, Think-J consists of two steps:
@@
2) Judgment Thinking Optimization. Due to the lack of high-quality critique annotation in preference datasets, we opt to optimize judgment thinking ability based on reinforcement learning (RL). Specifically, we adopted two methods: a) Critic-guided Offline Learning, leveraging an additional critic model to generate corresponding thinking traces based on provided judgment results, thus constructing positive and negative examples for offline RL. b) Rule-based Online Learning, defining rule-based rewards based on the correctness of judgment results, thus optimizing the thinking trace by online RL (Shao et al. 2024).
@@
    We propose to stimulate the judgment thinking ability of generative models with carefully curated data.
@@
    We propose to optimize the judgment thinking ability of generative models with reinforcement learning.
@@
    Our proposed Think-J significantly outperforms previous generative and classifier-based LLM-as-judges.
2 Background
After the emergence of LLMs, numerous efforts have been made to design a more effective method for LLM evaluation (Chang et al. 2023). One of the most scalable and effective methods is LLM-as-a-Judge (Li et al. 2023b; Zheng et al. 2023), namely utilizing proprietary LLMs, especially GPT4 (Achiam et al. 2023), to evaluate the LLM’s response. For example, AlpacaEval (Li et al. 2023b) used the win rate compared with baseline response determined by GPT-4 as the evaluation result. MT-Bench (Zheng et al. 2023) automatically scored the model’s answers using GPT-4 as the results. The GPT-4-based evaluator is proven to presents comparable or even better consistency compared with human.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, follow-up works suggest fine-tuning language models locally for evaluations, including JudgeLM (Zhu et al. 2023b), Auto-J (Li et al. 2023a), Prometheus (Kim et al. 2023), Prometheus-2 (Zhu et al. 2023a), OffsetBias (Park et al. 2024), etc. These work typically construct preference data with judgment annotations and then finetune open-sourced LLMs to generate the judgment. Despite these fine-tuned judge models all achieve comparable accuracy with proprietary models, the evaluation is mostly conducted on the in-domain testsets, and these works are verified with a low scalability on more general benchmarks (Huang et al. 2024).
Another group of work fine-tunes a classifier on preference data based on the Bradley-Terry model, which is more commonly used on reward modeling. This approach is simple yet effective, as demonstrated on RewardBench (Lambert et al. 2024) where most top-performing models are trained in a classification style (Liu et al. 2024a). However, this method does not fully leverage the generative capabilities of LLMs, and is unable to provide rationales for its judgments, which is crucial for scalable evaluation. While recent work has begun to leverage the generative abilities of LLMs to combine critiques for scalar reward prediction (Ke et al. 2024; Ye et al. 2024), these critiques are often distilled from stronger proprietary models and hardly influence the final prediction (Liu et al. 2025d). Effectively integrating the generative abilities of LLMs into evaluation remains an open challenge (Chen et al. 2025a; Whitehouse et al. 2025; Chen et al. 2025b; Wang et al. 2025a).
3 Methodology
@@ 3.1 Judgment Thinking Initialization
Recent studies have shown that LLMs inherently possess long chain-of-thought (CoT) reasoning capabilities, which can be activated with a small amount of data (Muennighoff et al. 2025; Ye et al. 2025; Liu et al. 2025a). In this work, we also curate high-quality preference data, LIMJ707, to initialize the thinking capability of the judge model. Specifically, LIMJ707 is selected based on three principles:
  • Accuracy: The judgment (preference) annotation should be correct. We leverage the high-quality preference data Skywork-Preference-v0.2111huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.2, which has been carefully validated to ensure accurate annotation.
@@
  • Diversity: The instruction should encompass various types to enhance judgment thinking capabilities in different aspects. We represent the instructions with an embedding model and then merge duplicate samples [^2] For more details please refer to Appendix B.1..
The data statistics during processing are shown in Table 1. Based on LIMJ707, we construct judgment thinking trace by Deepseek-R1. After that, the annotated samples are used to initialize the model with judgment thinking capability by Supervised Fine-tuning (Ouyang et al. 2022) [^3] Due to the overly long thinking trace generated by Deepseek-R1, we performed trace-clipping to reduce training overhead and improve efficiency. Please refer to Appendix B.1 for more details..
@@ 3.2 Judgment Thinking Optimization
To further enhance the alignment between the judge and human preference, the initialized thinking trace should be further optimized on preference data. However, preference data typically only includes binary labels without thinking trace annotations. Therefore, we propose judgment thinking optimization based on reinforcement learning (RL). Specifically, we propose two methods based on offline and online learning respectively, as shown in Figure 2.
Critic-guided Offline Learning
Offline RL methods represented by Direct Preference Optimization (DPO) (Rafailov et al. 2023) has been widely applied to LLM pipelines due to their efficiency and simplicity (Grattafiori et al. 2024). In this work, we also aim to optimize the judgment thinking ability based on offline learning.
Due to the lack of golden thinking annotation in preference dataset [^4] The rationale for fine-tuning an additional critic model over sampling to generate preference data is presented in Appendix A.3., we propose to train an additional critic model [^5] Notice the critic model here differs from the critic model in traditional RL which is used for advantage estimation., to help to construct training samples. Both the critic and the judge models are trained on the same data (i.e., LIMJ707), with the following distinctions:
@@
With the help of critic model, samples with thinking annotation are constructed based on the correctness of judgment result. Therefore, the judge model will be enhanced to generate more accurate thinking for better judgment.
Rule-based Online Learning
The recent success of R1-style methods have demonstrated the effectiveness of online RL using discrete, rule-based rewards (Shao et al. 2024). In this work, we also apply online rule-based RL approach to optimize the judgment thinking capability. More specifically, we mainly utilize the GRPO algorithm, with the optimization objective as follows:
@@
[…]
@@
$r_{\text{accuracy}}$ $=\begin{cases}1,&\text{if }\text{judgement}=\text{label}\\ 0,&\text{if }\text{judgement}\neq\text{label}\end{cases}$ (3)
$r_{\text{format}}$ $=\begin{cases}0,&\text{if }\text{format is right}\\ -0.5,&\text{if }\text{format is wrong}\end{cases}$ (4)
[…]
@@
where detailed weights for different rewards are presented in Section 5.2. Notice we incorporate a reward for assessing preference strength, which is defined as the degree to which the judge favors one response over another [^7] For example, a strength of 1 means the chosen response is only slightly better than the rejected, while a strength of 3 means the chosen is much better than the rejected. For more details about the criteria for the strength annotation, please refer to Appendix B.4.. The prompt template of judgment is also adjusted as:
@@
Preference strength helps to perceive the relative quality of response pairs, without uniformly providing the same reward for different pairs despite the quality gap. We employ a comparatively simple scale of reward scores of reward, as during actual training, we observed that the model tends to manipulate scores towards extreme values [^8] Please refer to Appendix A.1 for more details ..
While absolute score annotation for a given response is challenging, annotating relative preference strength is more readily accessible (Wang et al. 2024d, 2025b). Moreover, if absolute score annotation is absent in the preference data, we can firstly fine-tune a BT-classifier based on the data, then leverage the classifier to assess the relative strength between chosen and rejected responses (Wang et al. 2024b).
Model Sample Num RewardBench RMBench Auto-J
@@
Claude-3-5-Sonnet-20240620 96.4 74.0 81.6 84.7 84.2 68.9 70.7
@@
GPT-4o-2024-08-06 96.1 76.1 88.1 86.6 86.7 68.8 69.8
Gemini-1.5-Pro-0514 92.3 80.6 87.9 92.0 88.2 74.4 68.1
JudgeLM-33B (Zhu et al. 2023b) 100K 90.1 51.0 85.7 39.7 66.6 49.6 45.3
Prometheus-7b-v2.0 (Zhu et al. 2023a) 40K 83.9 49.2 72.8 72.0 69.5 52.4 63.1
Prometheus-8x7b-v2.0 (Zhu et al. 2023a) 40K 93.0 47.1 80.5 77.4 74.5 57.4 68.5
Llama-3-OffsetBias-8B (Park et al. 2024) 276K 92.5 80.3 86.8 76.4 84.0 66.0 68.7
CompassJudger-32B (Cao et al. 2024) 2041K 97.4 65.6 85.1 87.1 83.8 69.4 80.7
STE-Llama3.1-70B (Wang et al. 2024c) 20K 96.9 85.1 89.6 88.4 90.0 65.3 72.0
SynRM-Command-R-35B (Ye et al. 2024) 5K 97.5 76.8 88.5 86.3 87.3
CLoud-Llama3-70B (Ke et al. 2024) 350K 98.0 75.6 87.6 89.0 87.6
Think-J-Qwen-2.5-32B (Helpsteer2-Pref) 9.8K 96.7 83.2 90.1 92.0 90.5 79.8 75.8
Table 2: Experiment results of top LLM-Judges on RewardBench, RMBench, and Auto-J-test. In this table, we report the best performance achieved by various LLM-Judges, trained on different base models and datasets.
Model Method $H$-RLHF Helpsteer2-Pref
@@
Table 3: Experiment results of different LLM judge training methods on RewardBench. In this table, we report the performance of various methods trained on the same base models (Llama3-8B-Instruct and Qwen2.5-7B-Instruct) and datasets ($H$-RLHF and Helpsteer2-Pref) to enable a more direct comparison.
@@ 4.1 Set-up
We mainly conducted experiments on two popular open-sourced models with their instruction version: Qwen-2.5 (Qwen et al. 2025) and Llama-3 (Grattafiori et al. 2024).
We primarily conducted our optimization on two datasets: HelpSteer2-Pref999huggingface.co/datasets/nvidia/HelpSteer2 and HH-RLHF101010huggingface.co/datasets/Anthropic/hh-rlhf. For a fair comparison, LIMJ707 was mixed into all training sets. While larger preference datasets are available, we excluded them from our training because our primary objective was to explore the most effective method for training LLM judges.
We mainly compare Think-J with the following generative LLM-judge approaches:
  • Direct Prompt Leverage the LLM to directly generate judgment without fine-tuning.
  • SFT (w/o CoT) Train a generative model by supervised fine-tuning to perform judgment without thinking traces.
  • SFT (w/ CoT) Train a generative model on correct judgment thinking traces and results.
We also compare Think-J with the following classifier-based LLM-judge approaches:
  • BT Classifier Feed the instruction and responses into the model and added a classification head, training it according to Bradley-Terry model (Sun et al. 2025).
  • CLoud (Ke et al. 2024) First train the model to generate critiques, and then leverage the critiques as additional input to improve the BT Classifier.
  • SynRM (Ye et al. 2024) Leverage critiques generated by a proprietary model as additional input to improve the BT Classifier. We use Deepseek-R1 to generate the critiques.
@@
We mainly perform evaluation on RewardBench (Lambert et al. 2024). We also evaluate our model on RMBench (Liu et al. 2025b) and Auto-J-test (Li et al. 2023a).
We report the best results achieved by either offline or online learning for the main experiments by default.